Please show work!Problem-2. In the circuit shown below. (a) What is the total resistance as seen by the battery? (b) What is the current throu
Answer #1

3X 6 R23 2 2 (a) 3+6 R234 11 2+4= 6 R R2345 = 6x3 2 R 6+3 Reg= 3 +2= Jr (ط) lo I, - I - 2A W (c) p = VI = 20 lo X2= (d) V, 3x